OpenStudy (anonymous):

What is the sum? 7/3+(-3/8) a. 65/24 b. 47/24 c. -4/5 d. -5/4

OpenStudy (anonymous):

and can you show me how you got it?

OpenStudy (anonymous):

Ok so first find a common denominator: In this case it's 24 so what do I do to get 24? 7/3 times 8/8= 56/24 now -3/8 times 3/3= -9/24 add them together 56/24+-9/24 47/24

OpenStudy (anonymous):

so b